Northeast Baptist buys $500,000 of a particular item (at gross prices) from its major supplier, Cardinal Health, which offers NE Baptist terms of 3/20, net 60. Currently, the hospital is paying the supplier the full amount due on Day 60, but it is considering taking the discount, paying on Day 20 and replacing the trade credit with a bank loan that has a 12 percent rate.

a. What is the amount of free trade credit that Baptist obtains from Cardinal Health? (Assume 360 days per year throughout this problem.)

b. What is the amount of costly trade credit?

c. What is the approximate annual cost of the costly trade credit?

d. Should Baptist replace its trade credit with the bank loan? Explain your answer.

e. If the bank loan is used, how much of the trade credit should be replaced?
Answer #1

Answer (a):

Net Yearly purchase value = 500000 * 500000 * 3% = 500000 - 15000 = $485,000

Free trade credit is average sales for 20 days = 485000 * 20/360 = $26,944

Amount of free trade credit that Baptist obtains from Cardinal Health = $26,944

Answer (b):

Amount of costly trade credit is average sales for (60-20=) 40 days = 485000 * 40/360 = $53,889

Amount of costly trade credit = $53,889

Answer (c):

Annual cost of the costly trade credit = (Discount in %)/(100%- Discount in %) * 360 / (Payment Date-Discount Period)

= 3% / (100% - 3%) * 360/ (60 - 20)

= 27.84%

Annual cost of the costly trade credit = 27.84%

Answer (d):

Yes.

The bank loan rate is lower than the costly credit rate, so Baptist should replace the costly credit portion of the credit with a bank loan.

Annual cost of cost trade credit is = 27.84%

Annual cost of trade credit with a bank loan = 12%

So Baptist should take bank loan and pay on 20th Day (Discount day) and take discount.

Answer (c):

If the bank loan is used, cost trade credit should be replaced.

Hence:

If the bank loan is used, trade credit that should be replaced = $53,889